Analyze the annual reports of Flour Mills of Fiji (2020), RB Patel Ltd (2020) and Paradise Beverages Ltd (2019) from South Pacific Stock Exchange (SPSE) and answer the following questions. 

 

1.     Calculate the share value using Dividend Discount Model, assuming dividend will grow 3% for Flour Mills of Fiji, 6% for RB Patel Ltd and 9% for Paradise Beverages constantly for the companies on their respective base year. 

 

2.     Calculate the Earnings Per Share (EPS) and the Market Capitalization for each of the companies. 

 

3.     Comment on the share value of the companies given the Dividend Discount Model (Question 1) and EPS (Question 2).  


1
Expert's answer
2021-04-28T09:49:40-0400

Answers:

  1. Let the dividend be 9,10,11. Required rate of return 10

"P=\frac{D}{r-g}"

  • "P1=\frac{9}{0.1-0.03}=128.57"
  • "P2=\frac{10}{0.1-0.06}=250"
  • "P3=\frac{11}{0.1-0.09}=1100"


2. (i) Determination of Earnings Per Share (EPS)

"EPS=\frac{In-D}{Sa}"

  • "EPS1=\frac{12-9}{300}=0.01"
  • "EPS2=\frac{14-10}{200}=0.02"
  • "EPS3=\frac{15-11}{600}=0.007"


(ii) Determination of market capitalization (MC):

"MC=N\times P"

"MC1=300\times 128.57=38 571"

"MC2=200\times 250=50 000"

"MC3=600\times 1100=660 000"


3.

With the growth of dividends and the annual growth of interest on dividends, the share price also increases. As dividends , earnings, and share amounts rise, earnings per share fall